# What-if Analysis for Business Users: Current Practices and Future Opportunities

### v3

@@ 8.4. Limitations
We acknowledge the limitations of our studies. First, we are aware of the limited participant pool, representation of business users across four departments. Despite this, similar number of participants have been studied in previous visualization and HCI research studies (Kandel et al., 2012; Dimara et al., 2021; Alspaugh et al., 2019), and we chose few participants and covered multiple departments to ensure in-depth interviews and not have an extremely narrow collection of empirical data.
Second, our findings in the interview study are based on participants’ talk-through and self-reported experiences rather than actual workflows from their own use cases. Time constraints and the need for understanding their advanced analyses techniques comprehensively led us to choose verbal articulation over demonstration. We addressed this by asking them to talk us through a specific but common marketing use case that was familiar to participants. Additionally, in the follow-up task-based study, participant’s hands-on experience of using WIA techniques was captured. Nonetheless, we used only one use case for both parts of the study. However, this was necessary to minimize participant variance, facilitate fair analyses comprehension, and ensure participants’ comfort in sharing their analyses without compromising sensitive information.
For the similar reasons, we used the marketing-specific use case even for non-marketing participants. To accommodate for this limitation, we provided examples of how the MMM task would be relevant to other departments, such as how MMM could enhance sales efficiency through optimized channel management and promotional strategies for sales managers. Additionally, we encouraged participants from the non-marketing background to draw parallels of the techniques to their own use cases, enriching our findings with other contextualized examples.
@@ 9. Conclusion
In this paper, we conducted a two-part study with 22 business users from real-world enterprises. The first interview study aimed to understand the WIA techniques employed, tools used, and challenges faced in making data-driven decisions. Our findings show that business users independently perform a range of advanced analyses due to various logistical constraints, but often rely on rudimentary methods that are insufficient for effective decision-making. To further learn business users’ feedback on benefits and opportunities of enhancing advanced analysis techniques, we conducted a follow-up task-based study with the same participants. In this study, they used four representative WIA techniques identified in the first interview study, that we implemented in a visual analytics prototype as a probe for them to use hands-on to make data-driven decisions. Our findings showed that these techniques improved decision-making speed and confidence while also highlighting the need for better dataset preparation, risk assessment, and domain knowledge integration. Finally, based on these insights, we make design recommendations for future business analytics systems.
@@ Acknowledgements.
We thank Joe Hellerstein, Arjun Srinivasan, and the anonymous reviewers for their helpful feedback on the paper draft.

### v4

@@ 8.4. Limitations
We acknowledge the limitations of our studies. First, our participant pool of business professionals across four departments has a limited sample size per department. Despite this, a similar number of participants have been studied in previous visualization and HCI research studies (Kandel et al., 2012; Dimara et al., 2021; Alspaugh et al., 2019), and we chose few participants and covered multiple departments to ensure in-depth interviews and not have an extremely narrow collection of empirical data.
Second, our findings in the interview study are based on participants’ self-reported experiences rather than performed workflows from their own use cases. Time constraints and the need to understand their advanced analysis techniques comprehensively led us to choose verbal articulation over demonstration. We addressed this by asking them to talk us through a specific but common marketing use case that was familiar to participants. Additionally, in the follow-up task-based study, participants’ hands-on experience of using representative WIA techniques implemented in an interactive prototype was captured. We recognize that using such a prototype may introduce bias, where the specific visualization and interaction designs of the prototype can influence the results and the participants’ reactions. Future work needs to tease apart these potential confounding factors. Further, we acknowledge that this study focused solely on comparing the proposed representative techniques to participants’ current practices without evaluating them against other state-of-the-art tools. However, we believe that this step was essential to establish a baseline understanding of what business professionals perceive of advanced analytics techniques before introducing them to fully developed tools with extensive features, which is an exciting avenue for future work.
Further, we are aware that we used only one use case for both parts of the study. However, this was necessary to minimize participant variance, facilitate comprehension, and ensure participants’ comfort in sharing their analyses without compromising sensitive information. For similar reasons, we used the marketing-specific use case even for non-marketing participants. To accommodate for this limitation, we provided examples of how the MMM task would be relevant to other departments, such as how MMM could enhance sales efficiency through optimized channel management and promotional strategies for sales managers. Additionally, we encouraged participants from non-marketing backgrounds to draw parallels between the techniques and their own use cases, enriching our findings with other contextualized examples.
@@ 9. Conclusion
In this paper, we conducted a two-part study with 22 business professionals from real-world enterprises. The first interview study aimed to understand the techniques, tools, and challenges in utilizing WIA for data-driven decision-making. Our findings show that business professionals perform a range of advanced analyses without relying on data analysts due to various logistical constraints. However, they often use rudimentary methods that are insufficient for effective decision-making. We conduct a follow-up task-based study with the same participants to elicit business professionals’ feedback on the benefits and opportunities of enhancing advanced analysis techniques. In this follow-up study, the participants used four representative WIA techniques identified in the first interview study, which we implemented in a visual analytics prototype as a probe for them to use hands-on to make data-driven decisions. Our findings show that these techniques improved decision-making speed and confidence while also highlighting the need for better dataset preparation, risk assessment, and domain knowledge integration. Finally, we make design recommendations for future business analytics systems based on these insights.
@@ Acknowledgements.
We thank Sigma Computing Inc., where this work was initiated as an internship project. We also thank Joe Hellerstein, Arjun Srinivasan, and the anonymous reviewers for their helpful feedback on the paper draft.